Bramble Corporation sold $2,200,000, 7%, 5-year bonds on January 1, 2022. The bonds were dated January 1, 2022, and pay interest on January 1. Bramble Corporation uses the straight-line method to amortize bond premium or discount.

Prepare all the necessary journal entries to record the issuance of the bonds and bond interest expense for 2022, assuming that the bonds sold at 105

Jan 1.

Dec 31.
Answer #1
Debit Credit
Jan 1 Cash 2310000 =2200000*1.05
     Bonds payable 2200000
     Premium on Bonds payable 110000
Dec 31 Interest expense 132000
Premium on Bonds payable 22000 =110000/5
     Interest payable 154000 =2200000*7%